# Emerging Platforms Meet Emerging LLMs: A Year-Long Journey of Top-Down Development

### v2

@@ IV-C RQ2: Overhead of TapML
TABLE II: Testing time breakdown in TapML exemplified by deploying Llama2-7B on WebGPU with CUDA reference. “O(1)” stands for one-time effort.
In this section, we systematically examine the overhead incurred by TapML during development. For clarity, we exemplify the overhead analysis via a case study of deploying Llama2-7B on WebGPU by using CUDA as the reference platform, running on a workstation with an AMD Ryzen 9 5900X CPU and an NVIDIA RTX 3080 GPU. Specifically, Llama2-7B includes 32 decoder blocks and when being compiled to the general representation of our base framework, it leads to 456 operator kernels in 18 operator types. Consequently, to support Llama2-7B, we need to implement test 18 operator kernel functions and pass $456\times k$ operator-level tests, where $k$ is determined by the number and sizes of sequences to be completed.
Section IV-C depicts the testing time breakdown of supporting Llama-2 given 6 prefilling token and 1 decoding token. Specifically, testing the model includes two types of costs:
  1. 1.
    Rooting time: The time spent on running a model-wise validation on the target platform (i.e., WebGPU), including compiling the model into WebGPU code, initializing the model on WebGPU, and running the model on WebGPU with model-wise test inputs. These steps are used for end-to-end model validation which is indispensable regardless of the development methods employed. Specifically, the bottleneck comes from model compilation where compiling Llama2-7B fully on WebGPU results in 87% of the end-to-end testing time.
  2. 2.
    Overhead time: Besides routine end-to-end validation, we need to obtain a full CUDA compute graph for test carving and gradual offloading. While compiling, loading, and test-carving Llama2-7B on CUDA can take about 1 minute, these steps are one-time efforts and can facilitate the testing of all operators. For operator-wise validation, we need to use the universal runtime to run carved tests for each operator, whose cost is tiny, e.g., on average 0.1 seconds per operator.
@@
IV-D RQ3: Bugs in Emerging ML Systems
@@
Bugs by platform runtime. The emergence of new computing platforms often brings with it a degree of instability, including inconsistent or incomplete implementations to its specifications. For instance, we found one of our WebGPU programs runs flawlessly on Google Chrome but produces wrong results on the Technology Preview (Release 188) version of Safari on the same MacBook. Our diagnosis shows that such discrepancies can be attributed to the incomplete WebGPU API support on Safari, such as onSubmittedWorkDone, whose incompleteness leads to a synchronization bug in which unready buffers were accessed. Furthermore, the base framework by default reuses shared memory buffers of different data types, which is okay for CUDA but introduced Vulkan compilation failures as this feature is not permitted by Vulkan’s design. Additionally, WebGPU is subject to varying constraints across different devices. For example, when deploying Gemma on WebGPU of the MacOS system, we assumed the maximum number of threads to be the same as that of the native Metal runtime. However, this leads to GPUPipelineError because WebGPU at most allows for 256 threads on Apple Silicon which is much smaller than our assumed value. Similarly, on the PC end, WebGPU allows for 1024MB maximum storage buffer sizes, yet it is limited to 128MB for mobile devices that over-requesting buffer sizes leads to an error of “Invalid BindGroup”.
Bugs by base framework. Our base framework, i.e., TVM, is a compilation-based framework that compiles an ML model from its internal representation into target code. When extending our base framework, a range of issues surfaced, predominantly from compiler errors from both the base and extended parts of the framework. For instance, we observed that our generated Metal code does not perform checks for unaligned data, which leads to incorrect results due to accessing unknown data. Meanwhile, missing instruction support in code generation is a major error type discovered by our carved tests. For example, certain compilation failures were encountered when offloading a matrix multiplication operator to the Metal platform, because of unsupported SIMD instructions such as the selection instructions. Furthermore, we also encountered performance issues. For example, our Metal code-generation pipeline initially did not support warp-level instructions, such as simd_shuffle, and therefore vanilla instructions are used which leads to sub-optimal performance. Meanwhile, we also used to assume an inefficient setting of warp sizes on Apple Silicon devices, leading to undesired processing speed. All of these errors are quickly located thanks to the test carving and gradual offloading mechanism in TapML, which isolates the debugging scope to a single operator each time.
@@
IV-E RQ4: Lessons and Best Practices
@@
Best practice of gradual offloading: RWKV [31] is an innovative neural network architecture with RNN architecture and Transformer-level performance. Distinct from the prevailing Transformer architectures, RWKV introduces a unique and specialized scan kernel, namely the wkv kernel. This kernel does not lend itself to straightforward representation as a mere aggregation of conventional operators, which requires support from scratch to deploy on nascent platforms.
@@
V Related Work
V-A Emerging ML Models and Platforms
Machine Learning has demonstrated remarkable capabilities across a variety of tasks. Specifically, Large Language Models [1, 32, 33, 34] have shown proficiency in linguistic tasks by accurately predicting subsequent tokens. Meanwhile, diffusion models [35, 36, 37, 38, 39] are adept at producing realistic images and videos in response to textual prompts. However, deploying these models presents significant challenges. They not only require substantial computational resources but also necessitate repeated executions within a single request.
As these models increase in size and complexity, the desire to deploy them broadly conveniently, and powerfully persists. Major high-performance-computing corporations such as NVIDIA, AMD, Intel, and Google have released cutting-edge AI accelerators. Concurrently, startups like SambaNova, Cerebras, and Tenstorrent are introducing chips with innovative architectures. Additionally, mobile SoCs have incorporated specialized Neural Processing Units (NPUs), e.g., Apple’s Neural Engine and Qualcomm’s Hexagon Tensor Processor.
@@ V-A Emerging ML Models and Platforms
At the higher level, several recent frameworks have been developed to catch up with the rapid evolution of models. Examples include vLLM [40], Hugging Face Transformers [41], llama.cpp [42], TensorRT-LLM [43], and MediaPipe [29]. However, these frameworks often focus on major or specific platforms, e.g., CUDA for NVIDIA GPUs. TapML aims to close the gap by pioneering the support and optimization of underrepresented emerging platforms, to maximize the possible scenarios of ML applications. Meanwhile, instead of cherry-picking specific platforms, the top-down methodology in TapML unleashes the productivity of developing emerging ML systems, which allows us to have the most comprehensive yet timely support of emerging platforms in public.
V-B Testing and Debugging in ML Frameworks
Software testing techniques for ML frameworks and compilers have been well-studied in recent years to improve their reliability. Specifically, due to the huge complexity of large-scale systems, fuzzing [44] has been applied to generate random ML models as test programs, to exercise the compilation and execution pipeline of ML frameworks. To this end, recent work has been focusing on generating valid and diverse ML models [45, 46, 47, 48, 49, 17, 18] (e.g., PyTorch programs) or IRs [50, 51]. For example, NNSmith [17] applies a solver-aided approach to define operator constraints and generate valid graphs through SMT solving. Besides test model generation, recent work also improves the reliability of ML systems through translation validation [52] and discovering more oracles [53, 54, 55]. Rather than generating random models to test the whole system comprehensively, TapML focuses on testing a vertical path to deploying one model each time.
@@ V-B Testing and Debugging in ML Frameworks
Debugging techniques have been used in ML system development to accelerate troubleshooting. PolyGraphy [22] is a tool aiming to diagnose a small error-inducing subgraph out of an ONNX model [56] that fails on NVIDIA TensorRT [4], via test input reduction [57]. Specifically, PolyGraphy either performs bisection or linearly removes the top and bottom layers to get smaller sequence models to reproduce the error. However, such a test input reduction mechanism is neither optimal nor efficient, as it assumes the input model to be a sequence that can actually be a graph and requires significant attempts of recompilation and testing. TapML addresses the limitations by opening the black box of the DL framework to directly isolate errors in small minimal subgraphs through test carving [14].
VI Threats to Validity
External. The main concern of external validity is if TapML can be applied to develop emerging ML systems. To alleviate the concern, we built TapML and conducted the experiments by diversifying the experimental configurations, including the supported models, platforms, and involved developers. Specifically, we have successfully applied TapML to deploy 82 important models on 5 emerging and serious platforms successfully. Meanwhile, to our best effort, we employ as many as 11 developers involved in the development and study. The employed developers have different years of experience ranging from 0 to over 5 years and we have been practicing the methodology for over a year.
Internal. A potential threat to internal validity is that there may be faults in our implementation and experiments. We perform systematic unit testing and code reviews during our development to alleviate this concern.
VII Conclusion
@@
TapML aims to address the critical software development challenges posed by the rapid evolution of ML models and the emergence of new computing platforms. By adopting a top-down approach and utilizing the universal runtime, TapML streamlines the deployment process on diverse platforms, significantly enhancing developer productivity. This method, characterized by automated testing and a migration-based strategy, has proven to be effective in real-world applications, as evidenced by the successful deployment of 82 emerging models across 5 novel platforms. The comprehensive case studies derived from these deployments offer great insights and best practices for developing ML systems on emerging platforms. Ultimately, we show that the TapML approach accelerates the deployment process while ensuring the reliability and efficiency of emerging models, bridging the gap left by traditional methods and existing frameworks.

### v3

@@ 4.3. RQ2: Productivity Improvements
Table 5. Breakdown TapML testing time, exemplified by deploying Llama2-7B on WebGPU with CUDA reference.
4.4. RQ3: Overhead
We also systematically examine the testing overhead incurred by applying TapML in model deployment, compared to what is necessary. For clarity, we exemplify the overhead analysis using a case study of deploying Llama2-7B on WebGPU, using CUDA as the reference platform. The hardware setup is based on a workstation with an AMD Ryzen 9 5900X CPU and an NVIDIA RTX 3080 GPU. Specifically, Llama2-7B includes 32 decoder blocks; when compiled to the general representation of our base framework, it leads to 456 operator kernels in 18 operator types. Consequently, to support Llama2-7B, we need to implement test 18 operator kernel functions and pass $456\times k$ operator-level tests, where $k$ is the number of inference tokens determined by the end-to-end testing prompts.
Section 4.3 depicts the testing time breakdown of supporting Llama-2 given 6 prefilled token and 1 decoding token. Specifically, the model testing time is dominated by two parts:
  1. (1)
    Rooting time: The time for running a model-wise validation on the target platform (i.e., WebGPU), including compiling the model into WebGPU code (i.e., WGSL), initializing the model on WebGPU, and running the model on WebGPU with model-wise test inputs. These steps are used for end-to-end model validation, which is indispensable regardless of the development methods employed. Specifically, the bottleneck comes from model compilation where compiling Llama2-7B fully on WebGPU results in 87% of the end-to-end testing time.
  2. (2)
    Overhead time: Besides routine end-to-end validation, TapML additionally requires a full CUDA compute graph for test carving and gradual offloading. Compiling Llama2-7B on CUDA, loading its weights, and performing test carving are one-time efforts, which in total can cost less than one minute. Furthermore, after implementing each operator, we validate the operator implementation using the carved tests, taking 0.1 seconds on average, which will happen as many times as the number of operator types (i.e., 18 times).
@@ 4.4. RQ3: Overhead
4.5. RQ4: Encountered Bugs when Developing Emerging ML Systems
@@
Bugs from platform drivers. The emergence of new computing platforms often brings with it a degree of instability, including inconsistent or incomplete implementations to its specifications. For instance, our approach led to the discovery of a confirmed concurrency bug in Safari, a mature and major browser used by billions of Apple users. We found that one of our WebGPU programs runs flawlessly on Google Chrome but mysteriously produces wrong results on the Technology Preview (Release 188) version of Safari on the same MacBook. Our diagnosis shows that such discrepancies can be attributed to the incomplete WebGPU API support on Safari, such as onSubmittedWorkDone, whose incompleteness leads to a synchronization bug in which unready buffers were accessed. This interesting bug finding highlights the effectiveness of TapML in capturing challenging and subtle bugs in foundational software stacks before its major release, contributing to developer efficiency and software reliability both at once. Furthermore, the base framework by default reuses shared memory buffers of different data types, which is okay for CUDA but introduced Vulkan compilation failures as this feature is not permitted by Vulkan’s design. Additionally, WebGPU is subject to varying constraints across different devices. For example, when deploying Gemma on WebGPU of the MacOS system, we assumed the maximum number of threads to be the same as that of the native Metal runtime. However, this leads to GPUPipelineError because WebGPU at most allows for 256 threads on Apple Silicon, which is much smaller than our assumed value. Similarly, on the PC end, WebGPU allows for 1024MB maximum storage buffer sizes; however, it is limited to 128MB for mobile devices that over-requesting buffer sizes, leading to an error of “Invalid BindGroup”.
Bugs by base framework. Our base framework, i.e., TVM, is a compilation-based framework that compiles an ML model from its internal representation into target code. When extending our base framework, a range of issues surfaced, predominantly from compiler errors from both the base and extended parts of the framework. For instance, we observed that our generated Metal code does not perform checks for unaligned data, which leads to incorrect results due to accessing unknown data. Meanwhile, missing instruction support in code generation is a major error type discovered by our carved tests. For example, certain compilation failures were encountered when offloading a matrix multiplication operator to the Metal platform because of unsupported SIMD instructions, such as the selection instructions. Furthermore, we also encountered performance issues. For example, our Metal code-generation pipeline initially did not support warp-level instructions, such as simd_shuffle, and therefore, vanilla instructions are used, leading to sub-optimal performance. Meanwhile, we also used to assume an inefficient setting of warp sizes on Apple Silicon devices, leading to undesired processing speed. All of these errors are quickly located, thanks to the test carving and gradual offloading mechanism in TapML, which isolates the debugging scope to a single operator each time.
@@
4.6. RQ5: Lessons and Best Practices
@@
Best practice of gradual offloading: RWKV (Peng et al., 2023) is an innovative neural network architecture with RNN architecture and Transformer-level performance. Distinct from the prevailing Transformer architectures, RWKV introduces a unique and specialized scan kernel, namely the wkv kernel. This kernel does not lend itself to straightforward representation as a mere aggregation of conventional operators, which requires support from scratch to deploy on nascent platforms.
@@
5. Related Work
5.1. Emerging ML Models and Platforms
Machine Learning has demonstrated remarkable capabilities across a variety of tasks. Specifically, Large Language Models (Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2024; Gunasekar et al., 2023) have shown proficiency in linguistic tasks by accurately predicting subsequent tokens. Meanwhile, diffusion models (Saharia et al., 2022; Podell et al., 2023; Betker et al., 2023; Blattmann et al., 2023; OpenAI, 2024) are adept at producing realistic images and videos in response to textual prompts. However, deploying these models presents significant challenges. They not only require substantial computational resources but also necessitate repeated executions within a single request.
As these models increase in size and complexity, the desire to deploy them broadly, conveniently, and efficiently persists. Major high-performance computing corporations such as NVIDIA, AMD, Intel, and Google have released cutting-edge AI accelerators. Concurrently, startups like SambaNova, Cerebras, and Tenstorrent are introducing chips with innovative architectures. Additionally, mobile SoCs have incorporated specialized Neural Processing Units (NPUs), e.g., Apple’s Neural Engine and Qualcomm’s Hexagon Tensor Processor.
@@ 5.1. Emerging ML Models and Platforms
At the higher level, several recent frameworks have been developed to catch up with the rapid evolution of models. Examples include vLLM (Kwon et al., 2023), Hugging Face Transformers (Wolf et al., 2020), llama.cpp (Gerganov, 2023), TensorRT-LLM (NVIDIA, 2024), SGLang (Zheng et al., 2023), and MediaPipe (Lugaresi et al., 2019). However, these frameworks often focus on major or specific platforms, e.g., CUDA for NVIDIA GPUs. TapML aims to close the gap by initiating and accelerating the support of underrepresented emerging platforms, to diversify the possible scenarios of ML applications.
5.2. Testing and Debugging in ML Systems
Software testing techniques for ML frameworks and compilers have been well studied in recent years to improve their reliability. These approaches mainly focus on fuzzing-oriented (Miller et al., 1990) solutions to develop interesting test cases and extensive test oracles.
For diverse test generation, recent work has been focusing on generating valid and diverse ML models (Wei et al., 2022; Xie et al., 2022; Su et al., 2023; Wang et al., 2023b; Deng et al., 2023; Liu et al., 2023a, b) (e.g., PyTorch programs) or IRs (Lin et al., 2023; Wang et al., 2023a). For example, NNSmith (Liu et al., 2023a) applies a solver-aided approach to define operator constraints and generate valid graphs through SMT solving. TapML does not perform model generation since the model to test (i.e., the model to be deployed) is determined in model deployment.
Besides test model generation, developing test oracles is also crucial. The oracle behind TapML is differential testing (McKeeman, 1998; Evans and Savoia, 2007), which has often served as the de facto oracle in existing ML system fuzzers (Wei et al., 2022; Xie et al., 2022; Su et al., 2023; Wang et al., 2023b; Deng et al., 2023; Liu et al., 2023a, b). For example, given the same model and model inputs, we run the model computation over different settings in terms of platforms (e.g., CPU and GPU), optimization, etc. More recently, novel differential testing oracles have also been proposed to detect even more bugs. This includes applying translation validation (Bang et al., 2022) to test ML compiler passes, inferring output equivalence via API relationship (Deng et al., 2022), and expanding testing scenarios (Yang et al., 2023; Chen et al., 2023; Wang et al., 2022).
@@ 5.2. Testing and Debugging in ML Systems
Debugging techniques have been used in ML system development to accelerate troubleshooting. PolyGraphy (NVIDIA, [n. d.]) is a tool aiming to diagnose a small error-inducing subgraph out of an ONNX model (Bai et al., 2019) that fails on NVIDIA TensorRT (NVIDIA, 2022), via test input reduction (Zeller and Hildebrandt, 2002). Specifically, PolyGraphy either performs bisection or linearly removes the top and bottom layers to get smaller sequence models to reproduce the error. However, such a test input reduction mechanism is neither optimal nor efficient, as it assumes the input model to be a sequence that can actually be a graph and requires significant attempts of recompilation and testing. TapML addresses the limitations by directly isolating errors in small minimal subgraphs through a migration-based strategy to implement and test one operator each time.
Last, our paper is closely related to software migration (Fleurey et al., 2007). However, instead of migrating code (Nguyen et al., 2014; Zhong et al., 2010), TapML focuses on platform migration (Behrang and Orso, 2019, 2018; Qin et al., 2019) of ML models. Such migration requires correct and efficient model implementation using target-platform languages and runtime. Cross-platform migration is challenging; therefore, the methodology of TapML encourages making the process incremental and the result observable per step to minimize the scope of errors.
6. Conclusion
@@
TapML aims to address the critical software development challenges posed by the rapid evolution of ML models and the emergence of new computing platforms. By adopting a top-down approach and utilizing the universal runtime, TapML streamlines the deployment process on diverse platforms, significantly enhancing developer productivity. This method, characterized by automated testing and a migration-based strategy, has proven effective in our real-world practice, evidenced by the successful deployment of 105 emerging models across 5 novel platforms. The comprehensive case studies derived from these deployments offer great insights and best practices for developing ML systems on emerging platforms. Ultimately, we show that the TapML approach accelerates the deployment process while ensuring the quality of emerging models, bridging the gap left by traditional methods and existing frameworks.
Data Availability
The artifact of this paper is available at https://zenodo.org/records/15117702. The core implementation in TapML has been merged and maintained as part of the MLC-LLM project at https://github.com/mlc-ai/mlc-llm.
Acknowledgement
This project would have been impossible without the open-source efforts of the communities behind Apache TVM and MLC-LLM. Jiawei Liu is supported by a Ph.D. fellowship provided by Amazon.